OpenStudy (anonymous):

Write the sum or difference in the standard form a + bi. ( 3 + 3i) - ( -6 + i)

OpenStudy (anonymous):

$$(3+3i)-(-6+i)=3+3i+6-i=\dots$$

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (anonymous):

(3 + 3i) - (-6 + i) = 9 + 2i

OpenStudy (anonymous):

3 + 3i + 6 - i = 9 + 2i 3 + 6 = 9 3i - i = 2i 9 + 2i is your answer.

OpenStudy (anonymous):

Medal, please. If you need a explanation, ask.

OpenStudy (anonymous):

@jjdees Do you understand?

OpenStudy (anonymous):

yea sorry i didnt see it until now! thaank you!

OpenStudy (anonymous):

9 + 2i a + bi

OpenStudy (anonymous):

Welcome.